# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@ 2.5.6. Inter-Patch Dependencies for AI-Generated Image Detection by Chen: 2024 (IPD-Net)
Nataraj et al. (Nataraj et al., 2019) 97.70 63.15 53.75 92.50 51.1 54.7 57.1 63.06 55.85 65.65 65.80 60.50 70.70 70.55 71.00 70.25 69.60 69.90 67.55 66.86
Zhang et al. (Zhang et al., 2019) 49.90 99.90 50.50 49.90 50.30 99.70 50.10 50.00 48.00 50.60 50.10 50.90 50.40 50.40 50.30 51.70 51.40 50.40 50.00 55.45
Wang et al. (Wang et al., 2020b) 100.0 80.49 55.77 64.14 82.23 80.97 50.66 56.11 50.00 87.73 92.85 52.30 51.20 52.20 51.40 53.45 55.35 54.30 52.60 64.41
Chai et al. (Chai et al., 2020) 68.81 53.02 55.76 59.24 52.64 77.49 55.78 59.65 48.80 65.57 61.69 52.26 58.53 60.72 58.21 55.78 56.58 55.05 61.24 58.78
Ojha et al. (Ojha et al., 2023) 100.0 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.5 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
Wang et al. (Wang et al., 2023a) 100.0 67.73 64.78 83.08 65.30 100.0 94.75 57.62 60.96 62.36 62.31 83.20 82.70 84.05 84.25 87.10 90.80 90.25 58.75 77.89
Tan et al. (Tan et al., 2023) 99.90 85.10 83.00 94.80 72.50 99.60 56.40 47.80 41.10 50.60 50.70 74.20 94.20 95.90 95.00 87.20 90.80 89.80 88.40 _
Corvi et al. (Corvi et al., 2023) 100.00 92.00 96.90 99.40 94.80 99.50 54.10 90.60 55.50 100.00 100.00 53.90 58.00 61.10 57.50 56.90 59.60 58.80 71.70 _
Zhu et al. (Zhu et al., 2023) 99.00 99.50 99.30 99.05 99.00 96.75 88.20 63.50 67.50 93.90 98.75 98.70 98.80 98.60 98.75 98.75 98.75 98.75 98.45 94.42
Liu et al. (Liu et al., 2024b) 100.0 99.33 99.67 99.46 99.83 97.07 77.53 81.11 65.50 82.32 96.79 90.70 98.30 95.90 98.75 92.40 93.95 93.00 94.90 92.45
Cao et al. (Cao et al., 2024) 100.0 97.40 97.50 97.50 96.20 98.65 73.85 93.00 75.00 92.75 93.20 77.35 98.70 96.60 98.80 87.75 89.95 88.70 97.00 92.10
Liu et al. (Liu et al., 2024a) 99.90 99.30 99.50 97.20 99.40 99.80 93.20 _ _ _ _ 76.10 98.60 94.90 98.70 94.40 94.70 94.20 98.80 _
Koutlis et al. (Koutlis and Papadopoulos, 2025) 100.0 99.30 99.60 88.90 99.80 99.50 80.60 90.60 68.30 89.20 90.60 76.10 98.30 88.20 98.60 88.90 92.60 90.70 95.00 _
Table 10. Comprehensive Evaluation of methods on GAN and diffusion-generated images using the UnivFD dataset (Ojha et al., 2023). Results are reported in average precision (%).
@@
Nataraj et al. (Nataraj et al., 2019) 99.74 80.95 50.61 98.63 53.11 67.99 59.14 68.98 60.42 73.06 87.21 70.20 91.21 89.02 92.39 89.32 88.35 82.79 80.96 78.11
Zhang et al. (Zhang et al., 2019) 55.39 100.0 75.08 55.11 66.08 100.0 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
Wang et al. (Wang et al., 2020b) 100.0 96.36 85.34 98.10 98.48 96.97 60.33 82.95 54.22 99.61 99.81 69.93 66.17 67.68 66.13 71.18 76.37 72.13 67.66 80.50
Chai et al. (Chai et al., 2020) 68.44 55.59 64.37 64.10 58.74 84.48 59.92 72.08 47.63 73.05 68.38 58.98 77.05 76.87 76.35 75.97 77.41 74.68 71.91 68.74
Ojha et al. (Ojha et al., 2023) 100.0 99.46 99.59 97.24 99.98 99.60 82.45 61.32 79.02 96.72 99.00 87.77 99.14 92.15 99.17 94.74 95.34 94.57 97.15 93.38
Wang et al. (Wang et al., 2023a) 100.0 76.73 72.80 97.06 68.44 100.0 98.55 54.51 65.62 97.10 93.74 94.29 95.17 95.43 95.77 96.18 97.30 97.53 68.73 87.63
Corvi et al. (Corvi et al., 2023) 100.00 98.60 99.80 100.0 99.80 100.0 94.70 99.80 87.70 100.00 100.00 73.00 86.80 89.40 87.30 86.50 89.90 89.00 96.10 _
Zhu et al. (Zhu et al., 2023) 99.95 99.95 99.92 99.92 99.92 99.25 91.38 61.23 72.66 97.90 98.88 99.30 99.85 99.51 99.85 99.50 99.46 99.19 99.47 95.64
Liu et al. (Liu et al., 2024b) 100.0 99.85 99.88 99.69 100.0 99.68 87.38 88.26 84.48 98.82 99.84 93.39 99.81 96.80 99.88 98.71 98.84 98.60 98.81 96.99
Cao et al. (Cao et al., 2024) 100.0 99.96 99.89 99.73 99.93 100.0 88.38 97.12 89.22 98.82 99.98 95.31 99.86 99.14 99.90 97.20 97.99 98.02 99.65 97.90
Tan et al. (Tan et al., 2023) 100.0 94.00 90.70 99.90 79.30 100.0 67.90 _ _ _ _ 100.0 99.10 99.20 99.20 93.20 95.10 94.90 97.30 _
Liu et al. (Liu et al., 2024a) 100.0 100.0 100.0 99.80 100.0 100.0 98.00 _ _ _ _ 92.00 99.80 99.10 99.90 99.10 99.40 99.20 99.80 _
Koutlis et al. (Koutlis and Papadopoulos, 2025) 100.0 100.00 99.90 99.40 100.0 100.0 97.90 97.20 94.90 97.30 99.70 96.40 99.80 98.30 99.90 98.80 99.30 98.90 99.30 _
2.6. Commercial Detection Methods
@@
Watermarking by DeepMind (DeepMind, 2024) is a commercial technology that introduce watermarking to the generated image invisible to human eyes, that work as basis for the detection of the synthesized image properly.
3. Datasets
3.0.1. ForenSynths by Wang: 2020
@@
The dataset used by Wang et al. (Wang et al., 2020b) is publicly available on their GitHub page, as referenced in the main paper. The training set consists of 724,000 images, including 362,000 real images and 362,000 fake images generated by ProGAN (Karras, 2017), trained on the LSUN dataset (Yu et al., 2015) across 20 different object categories. The testing dataset, which includes images generated by various other GAN models, is also accessible on the GitHub page.
3.0.2. Artifact Dataset by Rahman: 2023
@@
Rahman et al. (Rahman et al., 2023) introduced the Artifact dataset, containing 2,496,738 images, including 964,989 real images and 1,531,749 synthetic images generated using 25 different models (both GANs and diffusion models). The dataset includes diverse object categories representative of general real-world content.
3.0.3. SynthBuster by Bammey: 2023
@@
Bammey et al. (Bammey, 2024) introduced the SynthBuster dataset, consisting of 9,000 synthetic images, with 1,000 images generated by each of nine different diffusion models, namely: DALL·E 2 (Ramesh et al., 2022), DALL·E 3 (Ramesh et al., 2022), Adobe Firefly (Adobe Inc., 2024), Midjourney v5 (MidJourney, 2024), Stable Diffusion (Rombach et al., 2022) 1.3, 1.4, 2, and XL, and GLIDE (Nichol et al., 2021). The dataset is publicly available, with access details provided in the main paper. For real images, the authors recommend using the RAISE-1k dataset (Dang-Nguyen et al., 2015), accessible via the same portal.
3.0.4. DiffusionForensics Dataset by Wang et al.: 2023
@@
Wang et al. (Wang et al., 2023a) introduced the DiffusionForensics dataset, specifically designed to evaluate forensic detectors for diffusion-generated images. The dataset includes images sourced from LSUN-Bedroom (Yu et al., 2015), ImageNet (Russakovsky et al., 2015), and CelebA-HQ (Karras, 2017), with the following distributions: - 42,000 synthetic images generated from LSUN-Bedroom, - 50,000 synthetic images generated from ImageNet, and - 42,000 synthetic face images generated using Stable Diffusion V2 (SD-V2), paired with 42,000 real images from CelebA-HQ.
3.0.5. UnivFD Dataset by Ojha et al.: 2023
@@
Ojha et al. (Ojha et al., 2023) expanded the ForenSynths dataset by integrating images generated from various models, including ProGAN (Karras, 2017), CycleGAN (Zhu et al., 2017), BigGAN (Brock, 2018b), StyleGAN (Karras et al., 2019), GauGAN (Park et al., 2019), StarGAN (Choi et al., 2018), Deepfakes (Rossler et al., 2019), SITD (Chen et al., 2018), SAN (Dai et al., 2019), CRN (Chen and Koltun, 2017), and IMLE (Li et al., 2019). In addition, diffusion-based models such as the Guided diffusion model (ADM) (Dhariwal and Nichol, 2021), LDM (Rombach et al., 2022), GLIDE (Nichol et al., 2021), and DALL-E (Ramesh et al., 2021) were included, increasing its diversity for forensic detection studies.
3.0.6. Community Forensics by Park: 2024
@@
Many existing datasets contain images generated by diverse generative models; however, they often lack diversity and generalization capability for many-to-many scenarios. These scenarios include testing images generated from multiple diffusion models against those from various GANs, as well as incorporating generative models such as VAEs. Recognizing the need for a large, diverse dataset capable of detecting synthetic images from unseen generative models, Park and Owens (Park and Owens, 2024) introduced the Community Forensics Dataset. This dataset comprises 2.4 million images, generated by 4,803 distinct generative models, alongside an equal number of real images, making it more diverse than previous datasets. To evaluate its effectiveness, the authors trained and tested state-of-the-art models on this dataset, utilizing a binary classification setup with pre-trained models ViT-S (Dosovitskiy, 2020) and ConvNeXt-S (Liu et al., 2022a) as backbone architectures. Unlike previous methods that freeze backbone layers, the authors fine-tuned the entire backbone end-to-end, achieving better performance compared to existing state-of-the-art methods on both the Community Forensics dataset and other publicly available benchmarks.
3.0.7. GenImage by Zhu: 2024

### v2

@@ 2.6.6 Inter-Patch Dependencies for AI-Generated Image Detection by Chen: 2024 (IPD-Net)
Wang et al. Wang et al. (2020b) 100.0 80.49 55.77 64.14 82.23 80.97 50.66 56.11 50.00 87.73 92.85 52.30 51.20 52.20 51.40 53.45 55.35 54.30 52.60 64.41
Chai et al. Chai et al. (2020) 68.81 53.02 55.76 59.24 52.64 77.49 55.78 59.65 48.80 65.57 61.69 52.26 58.53 60.72 58.21 55.78 56.58 55.05 61.24 58.78
Ojha et al. Ojha et al. (2023) 100.0 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.5 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
Wang et al. Wang et al. (2023a) 100.0 67.73 64.78 83.08 65.30 100.0 94.75 57.62 60.96 62.36 62.31 83.20 82.70 84.05 84.25 87.10 90.80 90.25 58.75 77.89
Tan et al. Tan et al. (2023) 99.90 85.10 83.00 94.80 72.50 99.60 56.40 47.80 41.10 50.60 50.70 74.20 94.20 95.90 95.00 87.20 90.80 89.80 88.40 _
Corvi et al. Corvi et al. (2023) 100.00 92.00 96.90 99.40 94.80 99.50 54.10 90.60 55.50 100.00 100.00 53.90 58.00 61.10 57.50 56.90 59.60 58.80 71.70 _
Zhu et al. Zhu et al. (2023) 99.00 99.50 99.30 99.05 99.00 96.75 88.20 63.50 67.50 93.90 98.75 98.70 98.80 98.60 98.75 98.75 98.75 98.75 98.45 94.42
Liu et al. Liu et al. (2024b) 100.0 99.33 99.67 99.46 99.83 97.07 77.53 81.11 65.50 82.32 96.79 90.70 98.30 95.90 98.75 92.40 93.95 93.00 94.90 92.45
Cao et al. Cao et al. (2024) 100.0 97.40 97.50 97.50 96.20 98.65 73.85 93.00 75.00 92.75 93.20 77.35 98.70 96.60 98.80 87.75 89.95 88.70 97.00 92.10
Liu et al. Liu et al. (2024a) 99.90 99.30 99.50 97.20 99.40 99.80 93.20 _ _ _ _ 76.10 98.60 94.90 98.70 94.40 94.70 94.20 98.80 _
Koutlis et al. Koutlis and Papadopoulos (2025) 100.0 99.30 99.60 88.90 99.80 99.50 80.60 90.60 68.30 89.20 90.60 76.10 98.30 88.20 98.60 88.90 92.60 90.70 95.00 _
Table 12: Comprehensive Evaluation of methods on GAN and diffusion-generated images using the UnivFD dataset Ojha et al. (2023). Results are reported in average precision (%).
@@
Wang et al. Wang et al. (2020b) 100.0 96.36 85.34 98.10 98.48 96.97 60.33 82.95 54.22 99.61 99.81 69.93 66.17 67.68 66.13 71.18 76.37 72.13 67.66 80.50
Chai et al. Chai et al. (2020) 68.44 55.59 64.37 64.10 58.74 84.48 59.92 72.08 47.63 73.05 68.38 58.98 77.05 76.87 76.35 75.97 77.41 74.68 71.91 68.74
Ojha et al. Ojha et al. (2023) 100.0 99.46 99.59 97.24 99.98 99.60 82.45 61.32 79.02 96.72 99.00 87.77 99.14 92.15 99.17 94.74 95.34 94.57 97.15 93.38
Wang et al. Wang et al. (2023a) 100.0 76.73 72.80 97.06 68.44 100.0 98.55 54.51 65.62 97.10 93.74 94.29 95.17 95.43 95.77 96.18 97.30 97.53 68.73 87.63
Corvi et al. Corvi et al. (2023) 100.00 98.60 99.80 100.0 99.80 100.0 94.70 99.80 87.70 100.00 100.00 73.00 86.80 89.40 87.30 86.50 89.90 89.00 96.10 _
Zhu et al. Zhu et al. (2023) 99.95 99.95 99.92 99.92 99.92 99.25 91.38 61.23 72.66 97.90 98.88 99.30 99.85 99.51 99.85 99.50 99.46 99.19 99.47 95.64
Liu et al. Liu et al. (2024b) 100.0 99.85 99.88 99.69 100.0 99.68 87.38 88.26 84.48 98.82 99.84 93.39 99.81 96.80 99.88 98.71 98.84 98.60 98.81 96.99
Cao et al. Cao et al. (2024) 100.0 99.96 99.89 99.73 99.93 100.0 88.38 97.12 89.22 98.82 99.98 95.31 99.86 99.14 99.90 97.20 97.99 98.02 99.65 97.90
Tan et al. Tan et al. (2023) 100.0 94.00 90.70 99.90 79.30 100.0 67.90 _ _ _ _ 100.0 99.10 99.20 99.20 93.20 95.10 94.90 97.30 _
Liu et al. Liu et al. (2024a) 100.0 100.0 100.0 99.80 100.0 100.0 98.00 _ _ _ _ 92.00 99.80 99.10 99.90 99.10 99.40 99.20 99.80 _
Koutlis et al. Koutlis and Papadopoulos (2025) 100.0 100.00 99.90 99.40 100.0 100.0 97.90 97.20 94.90 97.30 99.70 96.40 99.80 98.30 99.90 98.80 99.30 98.90 99.30 _
2.7 Commercial Detection Methods
@@
Watermarking by DeepMind DeepMind (2024) is a commercial technology that introduce watermarking to the generated image invisible to human eyes, that work as basis for the detection of the synthesized image properly.
3 Discussion and Limitation
Each category of forensic detection contributes distinct insights into identifying AI-generated imagery, forming a complementary ecosystem of methodologies. Spatial-domain and frequency-domain analyses remain the foundation of pixel- and signal-level forensic understanding. Spatial detectors such as DIRE Wang et al. (2023a) and SelfCon Boychev and Cholakov (2024) have moved beyond early CNN-based designs by integrating reconstruction and contrastive learning for improved generalization, whereas frequency-based frameworks like SAFE Li et al. (2024b) and AIDE Yan et al. (2024) extend robustness by coupling spectral cues with semantic embeddings. Together, these advances demonstrate that spatial–frequency fusion provides a more holistic view of generation artifacts, though both remain limited under heavy compression and geometric transformations. Fingerprint and patch-based methods address these weaknesses from orthogonal perspectives. Fingerprint-based detectors such as MaskSim Li et al. (2024a) and LGrad Tan et al. (2023) mitigate model dependence through spectral and gradient-informed fingerprints, achieving cross-generator consistency absent in earlier handcrafted designs. Patch-based frameworks, including IPD-Net Chen et al. (2024a) and PatchCraft Zhong et al. (2024), enhance sensitivity to localized texture inconsistencies while maintaining contextual awareness through inter-patch relationships. These directions collectively narrow the gap between local artifact discovery and global scene understanding. Successful yet, these methods require extensive training or finetuning, which is computationally expensive, as well as limiting in scope since they require retraining when new and advanced generative models enter the market. In recent years, these limitations have been addressed by training-free approaches such as RIGID He et al. (2024), HFI Choi et al. (2024), and ZED Cozzolino et al. (2024), which depart from traditional learning paradigms by requiring no training on AI-generated images, instead leveraging statistical stability, frequency aliasing, or entropy deviation. While these methods excel in scalability and generator-agnostic detection with low computational cost, they still lack interpretability and resilience to adversarial perturbations.
Even more recently, vision-language and reasoning-based multimodal models—spanning CLIP-based frameworks like HyperDet Cao et al. (2024) to reasoning-driven systems such as ForenX Tan et al. (2025) and ThinkFake Huang et al. (2025)—push the field toward semantically aligned and explainable forensics with reasoning. These approaches overcome the semantic blindness of classical detectors, offering text-grounded rationales for model decisions, though they rely heavily on computationally expensive foundation backbones and curated instruction data. Importantly, Fan et al. Fan et al. (2024) present detection that is fully dependent on LLMs without integrating understanding between text–vision data and still provide comparable results. While this is a great perspective on the usage of LLMs, it might suffer in generalizability or from hallucination, as LLMs may provide contextual reasoning with incorrect information during detection.
Finally, commercial detection frameworks unify several of these perspectives by prioritizing scalability and usability over transparency. They often integrate pretrained multimodal backbones and frequency cues for enterprise-level deployment, but their proprietary nature limits academic reproducibility and benchmarking.
For clarity, we present a comparative analysis of methods across six distinct categories, following a general review of their reported performances in the literature. The results are detailed in Table 3, Table 8, Table 10, and Table 5. Additionally, a comprehensive performance comparison across all six categories using the UnivFD dataset Ojha et al. (2023) is provided in Table 11 for classification accuracy and Table 12 for average precision. The evaluation results on the UnivFD dataset Ojha et al. (2023) are selected due to its inclusion of state-of-the-art detection methods, particularly those utilizing vision-language approaches, which have demonstrated strong forensic capabilities in synthetic image detection. The recent perspective of reasoning provided by Multimodal Large Language Models (MLLMs) helps in moving from a black box to a grey box understanding—explaining why content is generated for the general public. A limitation we faced for these methods’ quantitative comparative analysis is that they are recent and present new datasets, often compared with general MLLM models that were not designed for forensic detection; such comparisons differ from the scope of this study. Therefore, we only include the quantitative evaluation of ForenX Tan et al. (2025) by merging it with training-free methods in Table 5.
While previous methods from spatial-domain, frequency-domain, fingerprint-based, and patch-based categories provide good detection results on different datasets, they may suffer from generalizability issues and computational expense due to dependence on the training of generative methods. Training-free methods could provide generalizability and computational efficiency, yet they may fail when the generated images exhibit very high fidelity. MLLM (VLM) methods could offer strong detection performance with reasoning—aligning with recent trends in the AI field—but they may be limited by extensive computational requirements and may suffer from the inherent issue of LLM hallucination. A hybrid structure leveraging the cross-category functions of MLLM and training-free methods may provide a new perspective on robustness, reasoning, generalizability, and efficiency.
4 Datasets
4.0.1 ForenSynths by Wang: 2020
@@
The dataset used by Wang et al. Wang et al. (2020b) is publicly available on their GitHub page, as referenced in the main paper. The training set consists of 724,000 images, including 362,000 real images and 362,000 fake images generated by ProGAN Karras (2017), trained on the LSUN dataset Yu et al. (2015) across 20 different object categories. The testing dataset, which includes images generated by various other GAN models, is also accessible on the GitHub page.
4.0.2 Artifact Dataset by Rahman: 2023
@@
Rahman et al. Rahman et al. (2023) introduced the Artifact dataset, containing 2,496,738 images, including 964,989 real images and 1,531,749 synthetic images generated using 25 different models (both GANs and diffusion models). The dataset includes diverse object categories representative of general real-world content.
4.0.3 SynthBuster by Bammey: 2023
@@
Bammey et al. Bammey (2024) introduced the SynthBuster dataset, consisting of 9,000 synthetic images, with 1,000 images generated by each of nine different diffusion models, namely: DALL·E 2 Ramesh et al. (2022), DALL·E 3 Ramesh et al. (2022), Adobe Firefly Adobe Inc. (2024), Midjourney v5 MidJourney (2024), Stable Diffusion Rombach et al. (2022) 1.3, 1.4, 2, and XL, and GLIDE Nichol et al. (2021). The dataset is publicly available, with access details provided in the main paper. For real images, the authors recommend using the RAISE-1k dataset Dang-Nguyen et al. (2015), accessible via the same portal.
4.0.4 DiffusionForensics Dataset by Wang et al.: 2023
@@
Wang et al. Wang et al. (2023a) introduced the DiffusionForensics dataset, specifically designed to evaluate forensic detectors for diffusion-generated images. The dataset includes images sourced from LSUN-Bedroom Yu et al. (2015), ImageNet Russakovsky et al. (2015), and CelebA-HQ Karras (2017), with the following distributions: - 42,000 synthetic images generated from LSUN-Bedroom, - 50,000 synthetic images generated from ImageNet, and - 42,000 synthetic face images generated using Stable Diffusion V2 (SD-V2), paired with 42,000 real images from CelebA-HQ.
4.0.5 UnivFD Dataset by Ojha et al.: 2023
@@
Ojha et al. Ojha et al. (2023) expanded the ForenSynths dataset by integrating images generated from various models, including ProGAN Karras (2017), CycleGAN Zhu et al. (2017), BigGAN Brock (2018b), StyleGAN Karras et al. (2019), GauGAN Park et al. (2019), StarGAN Choi et al. (2018), Deepfakes Rossler et al. (2019), SITD Chen et al. (2018), SAN Dai et al. (2019), CRN Chen and Koltun (2017), and IMLE Li et al. (2019). In addition, diffusion-based models such as the Guided diffusion model (ADM) Dhariwal and Nichol (2021), LDM Rombach et al. (2022), GLIDE Nichol et al. (2021), and DALL-E Ramesh et al. (2021) were included, increasing its diversity for forensic detection studies.
4.0.6 Community Forensics by Park: 2024
@@
Many existing datasets contain images generated by diverse generative models; however, they often lack diversity and generalization capability for many-to-many scenarios. These scenarios include testing images generated from multiple diffusion models against those from various GANs, as well as incorporating generative models such as VAEs. Recognizing the need for a large, diverse dataset capable of detecting synthetic images from unseen generative models, Park and Owens Park and Owens (2024) introduced the Community Forensics Dataset. This dataset comprises 2.4 million images, generated by 4,803 distinct generative models, alongside an equal number of real images, making it more diverse than previous datasets. To evaluate its effectiveness, the authors trained and tested state-of-the-art models on this dataset, utilizing a binary classification setup with pre-trained models ViT-S Dosovitskiy (2020) and ConvNeXt-S Liu et al. (2022a) as backbone architectures. Unlike previous methods that freeze backbone layers, the authors fine-tuned the entire backbone end-to-end, achieving better performance compared to existing state-of-the-art methods on both the Community Forensics dataset and other publicly available benchmarks.
4.0.7 GenImage by Zhu: 2024